# Scalable Generalized Bayesian Online Neural Network Training for Sequential Decision Making

### v1

@@ 5.1 In-between uncertainty
In this experiment, we test the ability of HiLoFi to capture the in-between uncertainty [23] after a single pass of the data. We consider the one-dimensional dataset introduced in [61].
@@ Results.
Figure 1: In-between uncertainty induced by HiLoFi as we increase number of observations.
@@ 5.2 MNIST
We consider the standard MNIST classification problem, but in an online setting, where each image and its label is presented once, in a single pass over the data. For the classification problem, after predictions are made, the agent is revealed the true label of the data; we refer to this as the complete information case. We also consider the MNIST contextual bandit task introduced by [49], where the agent is presented with an image and must choose one of ten possible classes as an action, and then gets a binary feedback, based on whether its prediction was correct or not; we refer to this as the incomplete information case.
@@
For the bandit problem, we consider two exploration strategies: TS, which needs access to samples from an approximate posterior, and $\epsilon$-greedy, with $\epsilon=0.05$, which can leverage standard point estimation methods such as SGD or muon.
@@
We compare LLL, LRKF, LoFi, and HiLoFi. For LLL, we use the online last-layer variant from [13], without hyperparameter re-optimization at each timestep. In the $\epsilon$-greedy setting with LLL, we evaluate both online SGD with the AdamW optimiser [44] and the muon optimiser. Choice for hyperparameters are detailed in Appendix G.1.3.
Online learning results.
Figure 2 shows 5000-step rolling mean of the one-step-ahead classification outcome (one for correct, zero for incorrect) using the various learning algorithms. We observe that there is no clear difference between the methods. This shows that one can tackle this complete information problem without needing to model uncertainty. This is not the case in the incomplete information case that we study next, nor any of the other experiments below.
Figure 2: Rolling one-step-ahead accuracy for the MNIST dataset.
@@
Figure 3 shows the average cumulative reward over 10 runs of the contextual bandit version of MNIST. We observe that the top three performing methods are HiLoFi, LoLoFi, and LRKF. This outcome is expected because HiLoFi employs a full-rank covariance matrix in the last layer, which is better than a low-rank approximation of the covariance matrix associated with the last layer (LoLoFi), or a low-rank covariance to model all dependencies across layers (LRKF). For a detailed analysis of the trade-off between computation time and reward with error bars see Appendix G.1.3.
Figure 3: Cumulative reward for the bandit MNIST problem.
@@ 5.3 Recommender systems
We study the performance of the methods on the Kuairec dataset of [24], which is derived from a real-world, non-stationary, video recommender system. This dataset has been used to study non-stationary contextual bandits in [67].
@@
We group rows in the dataset (for a given user) in blocks of the 10 next videos that the user saw. For the features, we consider the $\log(1+x)$ transform of like count, share count, play count, comment count, complete play count, follow count, reply comment count, and download count. For the target variable, we use the $\log(1+x)$ transform of the watch ratio $[0,\infty)$.
@@
We perform sequential Bayesian inference using the relevant method, and then use TS to choose the action at each step (see Section 3.1). The choice of hyperparameters is in G.2.
@@ 5.4 Bayesian optimization
Following [7], we evaluate the competing methods on classical Bayesian optimization (BO) tasks, where the goal is to maximize an unknown function $g:[0,1]^{D_{\bm{x}}}\to\mathbb{R}$ using a surrogate model trained sequentially on past observations.
@@
We consider seven benchmark functions commonly used in BO [7]: Ackley (2D, 5D, 10D, 50D), Branin (2D), Hartmann (6D), and DrawNN (50D and 200D). These span a range of dimensionalities and multi-modalities, and are standard in evaluating global optimal performance. For all methods, except for DrawNN, we sample a function from the posterior predictive and evaluate it on a fixed set of candidate points generated using a Sobol sequence [55], in order to find its maximum. For DrawNN functions, we instead optimize the sampled function directly over the continuous domain $[0,1]$ using projected gradient descent, implemented with the Jaxopt library [5].
@@
We perform sequential Bayesian inference using the relevant method, and then choose the next query point at each step using TS (see Section 3.1). Results using expected improvement [30] are in Figure 11 in the Appendix.
@@ Results.
Figure 5 (left) shows the final best value on the vertical axis (higher is better), and the amount of run time on the horizontal axis (lower is better). The number of steps for each dataset are in the x-axis in Figure 10 in Appendix G.3. Our method is competitive with other methods. Figure 5 (right) shows the tradeoff between performance (measured by rank) and compute time. Our method sits on the Pareto frontier for a range of tradeoffs between time and performance; furthermore, across tasks, the performance and compute time of HiLoFi dominates (better performance and less compute time) those from competing methods except from VBLL (higher performance) and LRKF (lower compute time). Thus, HiLoFi offers strong tradeoffs between compute time and performance. For further analysis of these results see Appendix G.3.
@@ 6 Related work
Here we briefly discuss additional related work, beyond those in Section 1. We focus on deterministic methods for approximate online parameter inference. One line of work approximates the posterior using a diagonal plus low-rank (DLR) representation for the precision matrix, which allows for low-rank updates. This includes the L-RVGA method of [36], which applies a factor-analysis perspective to recursive updates, as well as SLANG [46] and LoFi [10]. Lastly, efficient low-rank Kalman filtering has been explored for state-space models in [53]. However, their method assumes that the dimension of the low-rank component is smaller than the dimension of the measurements, an assumption that is not required in our case. Going beyond low-rank approximations, the FOO-VB method [66] scales online BNN inference by considering a Kronecker product decomposition of the posterior (error-variance) covariance. In the offline setting, scaling BNNs using various low-rank assumptions has been studied in, e.g., [41, 59, 35].
Another class of methods draws on the lottery ticket hypothesis [38, 42], which suggests that sparse subspaces can yield performant models. Techniques such as the subspace neural bandit [17] and PULSE [9] restrict learning to a subset of weights represented as a linear subspace. However, these methods require offline learning of the projection matrix from the linear subspace to the space of weights before online inference begins.
@@ 7 Conclusion and future work
We introduced novel methods for efficient online training of neural networks. Our approach is based on linearization and optimal low-rank approximations of error covariance matrices. We proposed three variants: HiLoFi, LoLoFi, and LRKF. These variants offer a good trade-off between performance and runtime in static settings, and achieve highest performance and efficiency in high-dimensional, non-stationary scenarios.
One limitation of our methods is that it does not support fixed-lag smoothing, which can be important for sample efficiency in problems with low signal-to-noise ratio. In addition, our theory does not capture errors due to linearization. Additional limitations include sensitivity to outliers [1, 16, 37] and reliance on hyperparameters. In the future, we aim to extend our approach to fully-online reinforcement learning [20].
@@ Appendix A Notation
In what follows, ${\mathcal{M}_{m\times n}(\mathbb{R})}$ stands for the space of $m$-by-$n$ matrices with real values, ${\mathcal{M}^{++}_{m}(\mathbb{R})}$ is the space of positive definite (pd) matrices, and ${\mathcal{M}^{+}_{m}(\mathbb{R})}$ is the space of psd matrices. A vector ${\bm{a}}\in\mathbb{R}^{d}$ is a a column vector, $\mathbf{B}^{1/2}$ is the upper-triangular Cholesky decomposition of $\mathbf{B}\in{\mathcal{M}^{++}_{m}(\mathbb{R})}$, and $\mathbf{I}_{D}$ is the $D$-dimensional identity matrix with $D\in\mathbb{N}$.
The notation ${\cal Q}_{R}(\mathbf{B}_{1},\ldots,\mathbf{B}_{k})$ denotes the Cholesky factorization (in upper-triangular form) of the matrix $\sum_{k=1}^{K}\mathbf{B}_{k}^{\intercal}\,\mathbf{B}_{k}$, where each $\mathbf{B}_{k}$ is itself an upper-triangular Cholesky factor; see Appendix C.1 for details. The function ${\cal P}_{d}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})$ returns a $d\times D$ matrix which is the best rank-$d$ approximation (in rectangular form) of the matrix $\sum_{k=1}^{K}\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}$, where each $\mathbf{A}_{k}\in\mathbb{R}^{d_{k}\times D}$ is a rectangular matrix such that $\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}$ is symmetric and of rank, at most, $d_{k}$. That is, $\mathbf{C}^{\intercal}\,\mathbf{C}$ with $\mathbf{C}={\cal P}_{d}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})$ approximates the sum with a symmetric rank-$d$ matrix. Similarly, ${\cal P}_{d,+a}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})$ returns the best rank-$d$ approximation of $\sum_{k=1}^{K}\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}+a\,\mathbf{I}_{D}$, with $a>0$; see Appendix C.2 for details. The vector ${\bm{y}}_{1:k}$ is given by $[{\bm{y}}_{1}^{\intercal},\dots,{\bm{y}}_{k}^{\intercal}]^{\intercal}$, which is a vertically-stacked column vector. For vectors, $\|\cdot\|_{2}$ is the $L^{2}$ norm, and for matrices, $\|\cdot\|_{\rm F}$ is the Frobenius norm. Lastly, ${\cal N}({\bm{x}}\,|\,{\bm{m}},\mathbf{S})$ is a multivariate Gaussian density evaluated at ${\bm{x}}$ with mean ${\bm{m}}$ and covariance matrix $\mathbf{S}$.
@@ Appendix B Extended Kalman filtering for online learning
Here, we explain in more detail the background of the EKF for online learning presented in Section 3.2. To make this section self-contained, we repeat parts of the material introduced in that section. Let us assume that ${\bm{y}}_{t}={\boldsymbol{\ell}}_{t}^{\intercal}\,\phi({\bm{x}}_{t},{\bm{h}}_{t})+{\bm{e}}_{t}$, where ${\bm{e}}_{t}$ is a zero-mean random variable with observation covariance $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$. Given a starting vector ${\bm{\theta}}_{0}$, we assume that model parameters ${\bm{\theta}}_{t}\in\mathbb{R}^{D_{\bm{\theta}}}$ and observations ${\bm{y}}_{t}\in\mathbb{R}^{D_{\bm{y}}}$ evolve according to the state-space model
@@
We now consider a first-order approximation of ${f}$ around ${\bm{\theta}}_{t-1|t-1}$, that is,
${\bm{y}}_{t}\approx{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})+\tilde{\mathbf{J}}_{t}(\,{\bm{\theta}}_{t}-{\bm{\theta}}_{t-1|t-1})+{\bm{e}}_{t}\,,$ (7)
@@
$\operatorname*{arg\,max}_{\mathbf{A}\in{\mathcal{M}_{({D_{{\boldsymbol{\ell}}}}+{D_{{\bm{h}}}})\times(k\,{D_{\bm{y}}})}(\mathbb{R})}}\mathbb{E}\left[||{\bm{\theta}}_{t}-\mathbf{A}\,{\bm{y}}_{1:k}||_{2}^{2}\right],$ (8)
@@
$\boldsymbol{\Sigma}_{t|k}$ $:={\rm Var}({\bm{\theta}}_{t})-\mathbf{A}_{t|k}^{\star}\,{\rm Var}({\bm{y}}_{1:k})\,\mathbf{A}_{t|k}^{\star\intercal}.$ (11)
@@
$\boldsymbol{\epsilon}_{t}$ $={\bm{y}}_{t}-{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t}),$ (14)
@@
[…]
@@
$p({\bm{y}}_{t}\,|\,{\bm{y}}_{1:t-1})$ $=\int{\cal N}\left({\bm{y}}_{t}\,|\,{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})+\tilde{\mathbf{J}}_{t}(\,{\bm{\theta}}_{t}-{\bm{\theta}}_{t|t-1}),\mathbf{R}_{t}\right)p({\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t-1})\mathrm{d}{\bm{\theta}}_{t}$ (16)
$={\cal N}({\bm{y}}_{t}\,|\,\tilde{\mathbf{J}}_{t}\,{\bm{\theta}}_{t|t-1},\tilde{\mathbf{J}}_{t}\,\boldsymbol{\Sigma}_{t|t-1}\tilde{\mathbf{J}}_{t}^{\intercal}+\ \mathbf{R}_{t}).$

### v2

@@ 5.1 In-between uncertainty
In this experiment, we test the ability of HiLoFi to capture the in-between uncertainty [21] after a single pass of the data. This concept refers to how well a model captures uncertainty in regions of input space that lie between, or away from, observed data. Intuitively, we expect uncertainty to be higher in such unexplored areas. This behavior is important for Bayesian decision-making problems for effictive balancing of exploration and exploitation. We consider the one-dimensional dataset introduced in [58].
@@ Results.
Figure 1: In-between uncertainty induced by HiLoFi as a function of the processed observations.
@@ 5.2 MNIST
We consider the MNIST contextual bandit task introduced by [49], where the agent is presented with an image and must choose one of ten possible classes as an action, and then gets a binary feedback, based on whether its prediction was correct or not; we refer to this as the incomplete information case. An additional experiment on online classification is in Appendix G.1.1.
@@
We consider two exploration strategies: Classical TS and our proposed predictive-sampling (PS) as outlined in Table 1. Additional results comparison TS, PS, and $\epsilon$-greedy are in Appendix G.1.3.
@@
We compare LLL, LRKF, LoFi, and HiLoFi. For LLL, we use the online last-layer variant from [11], without hyperparameter re-optimization at each timestep. Choice for hyperparameters are detailed in Appendix G.1.3.
@@
Figure 2 shows the average cumulative reward over 10 runs of the contextual bandit version of MNIST. We observe that the top three performing methods are HiLoFi, LoLoFi, and LRKF. This outcome is expected because HiLoFi employs a full-rank covariance matrix in the last layer, which is better than a low-rank approximation of the covariance matrix associated with the last layer (LoLoFi), or a low-rank covariance to model all dependencies across layers (LRKF). For further results showing regret and comparison using $\varepsilon$-greedy, see Appendix G.1.3.
Figure 2: Cumulative average reward for the bandit MNIST problem.
Next, Table 2 reports the average cumulative reward and total runtime (over 10 trials) for each method, comparing posterior predictive sampling (pBayes) with Thompson sampling (TS).
Agent Reward (pBayes) Reward (TS) Time (pBayes) Time (TS)
HiLoFi 31,167.7 8.44
LoLoFi 30,038.5 3.67
LRKF 28,354.9 2.51
LoFi 23,184.5 23,304.3 4.76 38.90
LLL 13,922.3 7,201.9 1.36 191.55
Table 2: Performance comparison of Bayesian decision-making strategies: posterior predictive sampling (pBayes) vs. Thompson sampling (TS). Rewards are cumulative and averaged over 10 runs; time is in minutes.
Overall, pBayes is consistently faster across all methods. For the two agents where both approaches are applicable (LoFi and LLL), the average cumulative reward is nearly identical for LoFi, while pBayes clearly outperforms TS for LLL. However, TS incurs a much higher runtime due to the need to sample repeatedly from the high-dimensional posterior over model parameters, whereas pBayes only requires sampling in the comparatively low-dimensional outcome space (one dimension per arm/class). Finally, as expected, HiLoFi achieves the best performance, followed by LoLoFi and then LRKF. This performance hierarchy comes at a cost in runtime: HiLoFi is slower than LoLoFi, which in turn is slower than LRKF.
@@ 5.3 Recommender systems
We study the performance of the methods on the Kuairec dataset of [22], which is derived from a real-world, non-stationary, video recommender system. This dataset is used to study non-stationary contextual bandits in [65].
@@
We group rows in the dataset (for a given user) in blocks of the 10 next videos that the user saw. For the features, we consider the $\log(1+x)$ transform of like count, share count, play count, comment count, complete play count, follow count, reply comment count, and download count. For the target variable, we use the $\log(1+x)$ transform of the watch ratio $0,\infty)$.
@@
We perform sequential Bayesian inference using the relevant method, and then use TS to choose the action at each step (see Section 3.1). The choice of hyperparameters is in G.3.
@@ 5.4 Bayesian optimization
Following [5], we evaluate the competing methods on classical Bayesian optimization (BO) tasks, where the goal is to maximize an unknown function $g:[0,1]^{D_{\bm{x}}}\to\mathbb{R}$ with a surrogate model trained sequentially on past observations.
@@
We consider seven benchmark functions commonly used in BO [5]: Ackley (2D, 5D, 10D, 50D), Branin (2D), Hartmann (6D), and DrawNN (50D and 200D). These span a range of dimensionalities and multi-modalities, and are standard in evaluating global optimal performance. For all methods, except for DrawNN, we sample a function from the posterior predictive and evaluate it on a fixed set of candidate points generated using a Sobol sequence [54] to find its maximum. Instead, for Drawnn functions, we use projected gradient descent, implemented with the Jaxopt library [4], to optimize the sampled function directly over the continuous domain $[0,1]$.
@@
We perform sequential Bayesian inference using the relevant method, and then choose the next query point at each step using TS (see Section 3.1). Figure 12 in the Appendix shows the results using the expected improvement algorithm [29].
@@ Results.
Figure 4 (left) shows the final best value on the vertical axis (higher is better), and the amount of run time on the horizontal axis (lower is better). The number of steps for each dataset are in the x-axis in Figure 11 in Appendix G.4. Figure 4 (right) shows the tradeoff between performance (measured by rank) and compute time. Our method is on the Pareto frontier for a range of tradeoffs between time and performance; furthermore, across tasks, the performance and compute time of HiLoFi dominates (better performance and less compute time) those from competing methods except from VBLL (higher performance) and LRKF (lower compute time). Thus, HiLoFi offers strong tradeoffs between compute time and performance. For further analyses and results see Appendix G.4.
@@ 6 Related work
We briefly review deterministic approaches for approximate online parameter inference. Several methods maintain diagonal plus low-rank (DLR) structures for efficient covariance updates, including L-RVGA [36], SLANG [47], and LoFi [8]. Low-rank Kalman filtering has also been explored in state-space models [53], though with stronger restrictions on the rank relative to measurement dimension. Beyond low-rank, FOO-VB [64] exploits Kronecker-structured approximations, while offline methods explore similar ideas for scaling BNNs [41, 56, 34]. Another direction builds on the lottery ticket hypothesis [42, 38], restricting learning to sparse or low-dimensional subspaces. Examples include the subspace neural bandit [15] and PULSE [7], which pre-train projection matrices offline before online adaptation.
@@ 7 Conclusion and future work
We introduced a predictive-first approach for efficient online training of neural networks, based on linearization and structured low-rank approximations of error covariances. Our three variants achieve strong performance-runtime trade-offs, with HiLoFi particularly effective in high-dimensional, non-stationary settings. Limitations include the lack of fixed-lag smoothing, sensitivity to linearization and outliers [14, 37], and reliance on hyperparameters. Future work includes extending our approach to fully-online reinforcement learning [18].
@@ Appendix A Notation
Symbol Description
$\|\cdot\|_{2}$ $\ell_{2}$ norm.
$\|\cdot\|_{F}$ Frobenius norm.
${D_{\bm{y}}},{D_{\bm{x}}}\in\mathbb{N}$ Dimension of observations / inputs.
${D_{{\bm{h}}}},{D_{{\boldsymbol{\ell}}}}\in\mathbb{N}$ Number of parameters for the hidden layer / last layer.
${D_{\bm{\theta}}}={D_{{\bm{h}}}}+{D_{{\boldsymbol{\ell}}}}$ Total number of neural network parameters.
${\cal N}({\bm{x}}\,|\,{\bm{m}},\mathbf{S})$ Multivariate Gaussian density evaluated at $\bm{x}$} with mean $\bm{m}$} and covariance matrix $\mathbf{S}$.
$p_{\rm env}({\bm{y}}\,|\,{\bm{x}})$ Unknown data-generating process for $\bm{y}$} conditioned on $\bm{x}$}
$p_{{\bm{b}}}({\bm{y}}\,|\,{\bm{x}})$ Posterior predictive model for $\bm{y}$}, conditioned on $\bm{x}$}, and parameterized by $\bm{b}$}.
$\mathcal{M}_{m\times n}(\mathbb{R})$} Space of $m$-by-$n$ real-valued matrices.
${\mathcal{M}^{++}_{m}(\mathbb{R})}$ Space of $m$-dimensional positive definite (pd) matrices.
${\mathcal{M}^{+}_{m}(\mathbb{R})}$ Space of $m$-dimensional positive semidefinite (psd) matrices.
$\mathbf{I}_{D}$ $D$-dimensional identity matrix.
$\mathbf{S}^{1/2}$ Upper-triangular Cholesky decomposition of the pd matrix $\mathbf{S}$.
${\bm{x}}_{t}=({\bm{c}}_{t},{\bm{a}}_{t})$ Inputs with context ${\bm{c}}_{t}\in\mathbb{R}^{{D_{c}}}$ (possibly empty) and action ${\bm{a}}_{t}\in{\cal A}\subseteq\mathbb{R}^{D_{a}}$.
$y_{t}\in\mathbb{R}$ Scalar measurements, rewards, or observations obtained at timestep $t$.
${\bm{y}}_{t}\in\mathbb{R}^{D_{\bm{y}}}$ Measurements, rewards, or observations obtained at timestep $t$.
$\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$ Datapoint at time $t$.
${\bm{u}}_{1:t}=({\bm{u}}_{1},\ldots,{\bm{u}}_{t})$ Time-ordered collection of vectors ${\bm{u}}_{k}\in\mathbb{R}^{D_{{\bm{u}}}}$.
$\mathcal{D}_{1:t}=(\mathcal{D}_{1},\ldots,\mathcal{D}_{t})$ Dataset at time $t$.
${\boldsymbol{\ell}}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$ Last layer parameters.
${\bm{h}}\in\mathbb{R}^{D_{{\bm{h}}}}$ Hidden layer parameters.
${\bm{\theta}}=({\boldsymbol{\ell}},{\bm{h}})\in\mathbb{R}^{D_{\bm{\theta}}}$ Collection of all neural network parameters.
${f}({\bm{\theta}},{\bm{x}})={f}({\boldsymbol{\ell}},{\bm{h}},{\bm{x}})\in\mathbb{R}^{D_{\bm{y}}}$ Neural network output with parameters $\bm{\theta}$} and inputs $\bm{x}$}.
${\bm{e}}_{t}\in\mathbb{R}^{D_{\bm{y}}}$ Zero-mean noise with covariance $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$.
${\bm{u}}_{t}\in\mathbb{R}^{D_{\bm{\theta}}}$ Zero-mean dynamics noise with variance $\mathbf{Q}_{t}\in{\mathcal{M}^{+}_{{D_{\bm{\theta}}}}(\mathbb{R})}$.
$q_{{\boldsymbol{\ell}},t}\geq 0$ Last-layer dynamics.
$q_{{\bm{h}},t}\geq 0$ Hidden-layer dynamics.
${\bm{u}}_{t,{\boldsymbol{\ell}}}\in\mathbb{R}^{{D_{{\boldsymbol{\ell}}}}}$ Zero-mean dynamics noise of last layer parameters with variance $q_{{\boldsymbol{\ell}},t}\,\mathbf{I}_{D_{{\boldsymbol{\ell}}}}$.
${\bm{u}}_{t,{\bm{h}}}\in\mathbb{R}^{D_{{\bm{h}}}}$ Zero-mean dynamics noise of hidden-layer parameters with variance $q_{{\bm{h}},t}\,\mathbf{I}_{D_{{\bm{h}}}}$.
${\boldsymbol{\ell}}_{t|t}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$ Frequentist estimate for ${\boldsymbol{\ell}}_{t}$ given $\mathcal{D}_{1:t}$.
${\bm{h}}_{t|t}\in\mathbb{R}^{D_{{\bm{h}}}}$ Frequentist estimate for ${\bm{h}}_{t}$ given $\mathcal{D}_{1:t}$.
${\bm{\theta}}_{t|t}=({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t})\in\mathbb{R}^{D_{\bm{\theta}}}$ Frequentist estimate for ${\bm{\theta}}_{t}$ given $\mathcal{D}_{1:t}$.
$\boldsymbol{\Sigma}_{t}={\rm Var}({\bm{\theta}}_{t}-{\bm{\theta}}_{t|t})$ Error variance-covariance matrix.
$\tilde{\boldsymbol{\Sigma}}_{t}$ Algorithm-dependent surrogate covariance matrix to $\boldsymbol{\Sigma}_{t}$.
$\hat{\boldsymbol{\Sigma}}_{t}=\operatorname*{arg\,min}_{\boldsymbol{\Sigma}:{\rm rank}(\boldsymbol{\Sigma})=d}\left\|\tilde{\boldsymbol{\Sigma}}_{t|t}-\boldsymbol{\Sigma}\right\|_{\rm F}^{2}$ Best rank-$d$ approximation to the surrogate covariance matrix $\tilde{\boldsymbol{\Sigma}}_{t}$.
$\mathbf{C}_{t}\in{\mathcal{M}_{d\times D}(\mathbb{R})}$ Rectangular matrix $(d\times D)$ matrix such that $\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}=\hat{\boldsymbol{\Sigma}}_{t}$
$\tilde{\mathbf{L}}_{t+1}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t|t},\,{\bm{h}}_{t|t},\,{\bm{x}}_{t+1})$ Jacobian of neural network w.r.t parameters in last layer.
$\tilde{\mathbf{H}}_{t+1}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t|t},\,{\bm{h}}_{t|t},\,{\bm{x}}_{t+1})$ Jacobian of neural network w.r.t. parameter in hidden layers.
${\cal Q}_{R}(\mathbf{B}_{1},\ldots,\mathbf{B}_{k})$ Cholesky factorization of $\sum_{k=1}^{K}\mathbf{B}_{k}^{\intercal}\,\mathbf{B}_{k}$, where each $\mathbf{B}_{k}$ is an upper-triangular Cholesky factor (see Appendix C.1).
${\cal P}_{d}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})\in{\mathcal{M}_{d\times D}(\mathbb{R})}$ Best rank-$d$ approximation of $\sum_{k=1}^{K}\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}$, with $\mathbf{A}_{k}\in{\mathcal{M}_{d_{k}\times D}(\mathbb{R})}$, $d_{k}\leq D$.
${\cal P}_{d,+a}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})\in{\mathcal{M}_{d\times D}(\mathbb{R})}$ Best rank-$d$ approximation of $\sum_{k=1}^{K}\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}+a\,\mathbf{I}_{D}$, with $a>0$ (see Appendix C.2).
@@ Appendix B Extended Kalman filtering for online learning
Here, we explain in more detail the background of the EKF for online learning presented in Section 3.2. To make this section self-contained, we repeat parts of the material introduced in that section. Let us assume that ${\bm{y}}_{t}={\boldsymbol{\ell}}_{t}^{\intercal}\,\phi({\bm{x}}_{t},{\bm{h}}_{t})+{\bm{e}}_{t}$, where ${\bm{e}}_{t}$ is a zero-mean random variable with observation covariance $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$. Given a starting vector ${\bm{\theta}}_{0}$, we assume that model parameters ${\bm{\theta}}_{t}=({\boldsymbol{\ell}}_{t},{\bm{h}}_{t})\in\mathbb{R}^{D_{\bm{\theta}}}$ and observations ${\bm{y}}_{t}\in\mathbb{R}^{D_{\bm{y}}}$ evolve according to the state-space model
@@
We now consider a first-order approximation of f{f} around ${\bm{\theta}}_{t-1|t-1}$, that is,
${\bm{y}}_{t}\approx{f}({\bm{\theta}}_{t-1|t-1},\,{\bm{x}}_{t})+\tilde{\mathbf{J}}_{t}\,({\bm{\theta}}_{t}-{\bm{\theta}}_{t-1|t-1})+{\bm{e}}_{t}\,$, (12)
@@
[…]
@@
$\boldsymbol{\Sigma}_{t|k}$ $={\rm Var}({\bm{\theta}}_{t})-\mathbf{A}_{t|k}^{\star}\,{\rm Var}({\bm{y}}_{1:k})\,\mathbf{A}_{t|k}^{\star\intercal}.$ (16)
@@
$\boldsymbol{\epsilon}_{t}$ $={\bm{y}}_{t}-{f}({\bm{\theta}}_{t-1|t-1},{\bm{x}}_{t}),$ (19)
@@
Assume a Gaussian density for the initial parameters $p({\bm{\theta}}_{0})={\cal N}({\bm{\theta}}_{0}\,|\,{\bm{\theta}}_{0|0},\boldsymbol{\Sigma}_{0|0})$ and for the noise term $p({\bm{e}}_{t})={\cal N}({\bm{e}}_{t}\,|\,\boldsymbol{0},\,\mathbf{R}_{t})$. The BLUP ${\bm{\theta}}_{t|t}$ and the EVC $\boldsymbol{\Sigma}_{t|t}$ are the parameters of the Gaussian that characterize the posterior of model parameters ${\bm{\theta}}_{t}$ given ${\bm{y}}_{1:t}$, more precisely,
@@
$p({\bm{y}}_{t}\,|\,{\bm{y}}_{1:t-1})$ $=\int{\cal N}\left({\bm{y}}_{t}\,|\,{f}({\bm{\theta}}_{t-1|t-1},\,{\bm{x}}_{t})+\tilde{\mathbf{J}}_{t}(\,{\bm{\theta}}_{t}-{\bm{\theta}}_{t|t-1}),\mathbf{R}_{t}\right)p({\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t-1})\mathrm{d}{\bm{\theta}}_{t}$ (21)
$={\cal N}({\bm{y}}_{t}\,|\,{f}({\bm{\theta}}_{t-1|t-1},{\bm{x}}_{t}),\;\;\tilde{\mathbf{J}}_{t}\,\boldsymbol{\Sigma}_{t|t-1}\tilde{\mathbf{J}}_{t}^{\intercal}+\ \mathbf{R}_{t}).$